Hi, I want to make a planter box for my garden. I know someone with 20 pieces of 2" by 6" by 10'. If I want my box 10' long by 3' wide and 1'tall how many pieces should I get? I cannot get it!
Thank you, I hope you can help asap.

Answer by checkley77(12844)

You will need 6 pieces of 10'for the bottom
6*6"=36"=3' wide bottom
You'll need 4 10'pieces for the 2 sides
2*6"=12"=1' for the one side
2*6"=12"=1' for the other side.
You'll need 2 3' pieces for each end
2*3'=6' for one end
2*3'=6' for the second end
All together you'll need:
6+4+2=12 10' pieces.